Subject: 3.5-m Users Committee minutes 6/12/06

From: Bruce Gillespie

Submitted: Tue, 13 Jun 2006 12:56:48 -0600

Message number: 970 (previous: 969, next: 971 up: Index)

                         APO 3.5-m Users Committee Phonecon, 6/12/06

Attending:  Suzanne Hawley, Bruce Gillespie, Michael Strauss, John  
Wilson, John Bally, Don York, Karl Glazebrook, Al Harper, Russet  

Absent: Jon Holtzman, Remy Indebetouw, Bruce Balick

Minutes taken by Bruce Gillespie


User feedback, comments from institutional representatives:

Princeton (Michael Strauss):  Nothing much to report, everything is  

Colorado (John Bally):  Nothing new to report, except that CU staff  
expecting to start commissioning of NIC-FPS Fabry-Perot etalon soon.

NMSU:  No report.

Johns Hopkins (Karl Glazebrook):  Nothing to report.

Chicago: (Al Harper):  No complaints.

UWashington (Suzanne Hawley): Several UW grad students who were heavy  
users of the 3.5m are defending soon, so we are training and  
expecting use by some new, younger students starting this summer.

Univ. of Virginia: No report.


Discussion of telescope/instruments report:

BruceG summarized highlights from the report (see below).


      3.5-m Telescope, Instruments, and CIF Projects Highlights,  
4/28/06 through 6/08/06

                                  Bruce Gillespie

0) Overview

Highlights include continued dry, less dusty, but cloudier conditions  
at the site, and smooth operations except for a spate of various  
computer and network communication problems; successful installation  
of the new azimuth drive boxes; good progress on the new top-end, and  
not-so-good progress on the DIS-upgrade projects; and release of a  
new version of TUI.

1) Telescope

The 3.5-m telescope and its supporting systems were fairly robust  
during this period, and observing was done nearly any time the  
weather allowed.  Part of one night was lost to a problem with the  
serial communication link to the altitude drive controller, which was  
quickly diagnosed and repaired.  Not so quickly repaired was a fault  
with our 20 Mb/s microwave link to the internet.  It turned out to be  
caused by a flakey transceiver at the far end (Tularosa) of the link,  
which was replaced after taking nearly two weeks to determine that  
the problem wasn't just some kind of interference.  The ISP has since  
acquired test equipment and spare parts to fix these kinds of  
problems more quickly.  In the interim, we were able to revert to a  
backup T1 circuit that we had kept for such a contingency, which  
slowed remote operations somewhat but was generally usable for remote  
observing.  We also experienced a rare failure of the Telescope  
Control Computer.  We were able to configure the TCC spare to take  
over running the telescope within a day, but we learned some things  
about how hard it is to keep the TCC spare in readiness.  In  
addition, a GPS unit, which provides time to the systems, went down  
and we reverted to the spare.  This cost a little observing time, and  
an improved communication box to the GPS has been ordered to make our  
GPS system more solid.  Lastly, we continued to see occurrences of  
NA2 rotator overcurrenting.  Some adjustments were made to rotator  
and instrument balance, and the problem seems to be somewhat better.   
However, we know the rotator mechanics are marginal, and a study is  
underway for a future upgrade to the rotator.

2) Instruments

The performance of all the instruments was generally good.  After  
reports of a return of the DIS "blue-haze" scattered light problem,  
the blue dewar was warmed and pumped, after which the scattered light  
appeared to be gone.

3) Engineering and CIF projects

TUI was upgraded to version 1.2. The RAID s/w backup system was  
purchased and assembled. A large-screen weather display was installed  
in the control room.  The new az drive boxes were successfully  
installed and tested.  We completed a long-term maintenance project  
to fix various enclosure shutter problems.  Some guiding problems  
with the newer versions of TUI were seen; fixes were incorporated and  
tested in TUI 1.2.  DIS upgrades were replanned--the new blue-medium  
grating will be installed in June, but the other upgrades are being  
deferred due to delays in procurements of components (new red  
detector, other new gratings, blue prism).  The top-end project is  
going well and is scheduled for installation on the telescope  
beginning 28 August.

4) Miscellaneous

Fire risk in the Lincoln Nat'l forest remains high--the forest is now  
closed to the public, but APO remains open to visitors and users. The  
APOLLO project made a major breakthrough in technique that showed  
immediate promise in enabling reliable acquisition of lunar-return  


DIS upgrade plans - Hawley:

Suzanne reported that the new blue-medium grating has been received,  
characterized at JHU, and will be installed in DIS this coming week.   
Once we confirm its performance in situ, we plan to make it the  
default for blue low/medium spectra.  The other two new gratings have  
not yet been delivered to JHU, and their testing and installation  
will have to wait.  The other DIS upgrade items previously planned  
for the end of June have been postponed, due to vendor delays in  
delivery of the new blue-prism and the new red detector.


CorMASS and NIC-FPS grism - Hawley:

Suzanne pointed out that we have little-used low-resolution NIR  
spectrograph capabilities in both CorMASS and the NIC-FPS grism.   
CorMASS is a well-documented echelle spectrograph for point sources,  
and has an easily used data reduction package supported by UVa.  The  
NIC-FPS grism is relatively new, and once characterized, should be  
useful for extended sources.  It is unclear why APO users are not  
requesting to use these capabilities, so Suzanne suggested that we  
should find a way to promote their use more energetically.


Q3 schedule overview - Hawley:

The new Q3 schedule is out.  Russet pointed out that the new Q3  
schedule has Princeton and JHU undersubscribed, and UChicago  
oversubscribed.  Michael said that most of the Princeton users have  
mostly extragalactic programs, for which this isn't the season.  Karl  
said there were administrative problems getting the JHU requests in,  
since there is a new JHU scheduler; Suzanne said that there have  
since been requests from JHU users for the open time in the  
schedule.  Al mentioned that the UC demand for time has gone up,  
because there are new users at Chicago.  Suzanne also pointed out  
that the top-end installation shutdown starts at the end of August  
and runs about three weeks, impinging slightly on plans to use the  
telescope in September for SDSS-II SNe follow-up.


Post-2008 ARC Call for Proposals - Hawley, Gillespie:

A Call for Proposals was sent to the 3.5-m and SDSS mailing lists at  
the end of May, with a July 17 deadline, to solicit proposals for  
future use of the 2.5-m telescope after SDSS-II ends in mid-2008.   
The proposal call includes the possibility of collaboration with the  
3.5-m.  All interested users are encouraged to participate in the  
proposal process.  Copies of the CfP, and related information, may be  
obtained from Bruce Gillespie (  Suzanne and  
Don York pointed out that this CfP is the next step in the process  
begun with the ARC Futures committee document, now in near-final  
draft form.  Users Committee and Futures Committee representatives at  
each institution have copies of the draft and should be actively  
soliciting comments.  The final report is anticipated to be complete  
in the next few weeks.


Announcements - Hawley:

Suzanne noted that Gabrelle Saurage is starting work as our new  
Observing Specialist.  Also, Suzanne reiterated our recent policy  
announcement on science data storage at APO:  Starting next month, we  
are going to delete science data from the APO disks from Q3 2005 and  
earlier.  At the beginning of subsequent quarters, the data from the  
same quarter of the previous year will be erased.


Previous Meetings Action Items report:

[open from previous months]:

ACTION:  BruceG will ask Russell what the pros and cons are of  
keeping TUI compatible with older versions of operating systems and  
libraries.  STATUS:  Discussed with RO, near to final recommendation.

ACTION:  Users Committee members should poll their users to ask if  
there are compelling reasons to keep TUI operable without using the  
latest version of operating systems.  STATUS:  Now closed; little  
feedback so far.

ACTION:  BruceG to discuss issues of computer and network security  
with Fritz Stauffer.  STATUS:  Working group (site staff and users)  
formed to frame requirements.

ACTION:  UC members should poll their constituents about ToO  
observations on the telescope.  Do they envision doing this type of  
science? Would expanding ToO opportunities be useful to them?  How do  
they feel about being asked to voluntarily give up time, and to  
having their time pre-empted (perhaps with later payback)?  STATUS:   
Closed, see minutes last meeting.

ACTION:  BruceG will ensure that the orientation checklist is up to  
date and available on-line, so that prospective new users can review  
it prior to their orientation site visit.  STATUS:  Open, plan to  
complete in early June.

ACTION:  Jon Holtzman to talk to Fritz and BruceG about installing a  
RAID backup at NMSU.  STATUS:  Backup computer assembled and tested  
at APO.  Will shortly copy the backup data to the spare.  System to  
be installed at NMSU in next few weeks.

[new actions from last month]:

ACTION: Suzanne asked each Users Committee member to send her a  
paragraph or two in the next two days with an expression of interest  
and/or questions regarding this proposed NIR instrument.
STATUS:  Some questions and supportive comments received, sent to  
instrument PI for consideration.

ACTION:  BruceG to forward Futures documents to Users Committee  
members, who then help the institutional Futures Committee member  
share them with their faculties and gather feedback.
STATUS:  Draft sent to committee members.


Next Users Committee phonecon will be on Monday, 17 July, at 8:30 AM  
Pacific Time.  Agenda and other materials will be sent to the  
committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 970 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to